14 J of heat are removed from a gas sample while it is being compressed by a piston that does 30 J of work.

What is the change in the thermal energy of the gas?
Answer #1

According to the first law of thermodynamics we know that ,

Q = U + W.

Here ,W = -30 J

Q = -14 J

So, U = 30-14 = 16 J

Change in thermal energy = 16 J